calculate the [H+] in a solution that has a pH of 10.08.
a) 3.9M
b) 10.1M
c) 1.2 x 10^-4 M
d) 8.3 x 10^-11 M
e) none of these
Answer #1

pH is defined as negative logarithm of H+ concentration.

pH=-log([H+])

Here pH=10.08

So -log(H+)=10.08

log(H+)=-10.08

Taking antilogarithm on both sides

[H+] =10-10.08 =8.3×10-11M

So answer is d)8.3×10-11M